# Dynamically localized but deconfined excitations in strained classical spin ice

### v1

@@ B.1 Phase transition at $\delta>1$
B.2 Scaling collapse for type 1A monopoles
@@
The correlation time of the noise is expected to be inversely proportional to the density of the dynamically active monopoles, in this case, the type 1A monopoles. We have implicitly made a scaling collapse in panel (a) of Fig. 8 for $-1<\delta\leq 0$, since we chose values of $T$ and $\delta$ which keep $(2-2|\delta|)J/T$ constant. The density of type 1A monopoles goes as $\exp{\left((2-2|\delta|)J/T\right)}$ ) for negative $\delta$. was sidetracked by other projects…
@@
from Eq. (missing) 8. Indeed, we find a perfect scaling collapse for ${\rm PSD}_{[111]}$ curves at different $T$ and $\delta$ by plotting ${\rm PSD}_{[111]}/\kappa_{1}$ versus $\kappa_{1}\omega$. This is shown in Fig. 11 for the regime $10\leq\delta<$ 1.
@@ B.3 Further MSD and PSD results
The PSDs shown in Fig. 13 are measured along two orthogonal directions, providing a different perspective from the sublattice resolved PSD curves presented in the main text. The PSD along $[111]$ has a mixed contribution from monopole movement through $\mathcal{S}_{[111]}$ and $S_{K}$ as spins belonging to $S_{K}$ have non-zero components along the $[111]$ direction, while the PSD along $[1\bar{1}0]$ has contributions from $S_{K}$ spins only. The anisotropic diffusion of type 1A monopoles for $\delta<0$ is manifested due to dynamical constraints imposed in directions other than $[111]$, whereas signals of dynamical localization appear in the suppression of $[111]$ directional PSDs for type 1B monopoles at $\delta>0$.

### v2

@@ C.1 Phase transition at $\delta>1$
C.2 Scaling collapse for type 1A monopoles
@@
The correlation time of the noise is expected to be inversely proportional to the density of the dynamically active monopoles, in this case, the type 1A monopoles. We have implicitly made a scaling collapse in panel (a) of Fig. 8 for $-1<\delta\leq 0$, since we chose values of $T$ and $\delta$ which keep $(2-2|\delta|)J/T$ constant. The density of type 1A monopoles goes as $\exp\left((2-2|\delta|)J/T\right)$ for negative $\delta$.
@@
from Eq. (missing) 9. Indeed, we find a perfect scaling collapse for ${\rm PSD}_{[111]}$ curves at different $T$ and $\delta$ by plotting ${\rm PSD}_{[111]}/\kappa_{1}$ versus $\kappa_{1}\omega$. This is shown in Fig. 11 for the regime $10\leq\delta<$ 1.
@@ C.3 Further MSD and PSD results
The PSDs shown in Fig. 13 are measured along two orthogonal directions, providing a different perspective from the sublattice resolved PSD curves presented in the main text. The PSD along $[111]$ has a mixed contribution from monopole movement through $\mathcal{S}_{[111]}$ and $\mathcal{S}_{k}$ as spins belonging to $\mathcal{S}_{k}$ have nonzero components along the $[111]$ direction, while the PSD along $[1\bar{1}0]$ has contributions from $\mathcal{S}_{k}$ spins only. The anisotropic diffusion of type 1A monopoles for $\delta<0$ is manifested due to dynamical constraints imposed in directions other than $[111]$, whereas signals of dynamical localization appear in the suppression of $[111]$ directional PSDs for type 1B monopoles at $\delta>0$.